OpenStudy (anonymous):

Solve: 5(x + 6) less than or greater to 3x - 4 x greater than or equal to -17 x greater than or equal to 17 x less than or greater to -17 x less than or greater to -5

OpenStudy (anonymous):

Did you mean 5(x+6) less or EQUAL to 3x-4?

OpenStudy (anonymous):

Yes, sorry.

OpenStudy (anonymous):

Well first step is to distribute the 5 over (x+6)

OpenStudy (anonymous):

so, 5x+30

OpenStudy (anonymous):

yep ... now move all of the x terms to one side.

OpenStudy (anonymous):

2x+30 less or EQUAL to -4?

OpenStudy (anonymous):

yep now get the x term by itself...

OpenStudy (anonymous):

2x less or EQUAL to -34?

OpenStudy (anonymous):

yep now get it to be just x instead of 2x.

OpenStudy (anonymous):

Then do the math? and I get c? x less than or greater to -17

OpenStudy (anonymous):

Less than or equal ... yes.

OpenStudy (anonymous):

nice job!

OpenStudy (anonymous):

Thanks alot.

OpenStudy (anonymous):

5(x + 6) <= 3x - 4 5x + 30 <= 3x - 4 5x - 3x <= -30-4 2x <= -34 x <= - 17 you got it :)

OpenStudy (anonymous):

Thanks as well kellie

OpenStudy (anonymous):

lol...I didn't do anything...medal should go to Paynesdad